1. Two numebers have a sum of 26. Three times the first less twice the second is 3. Find the numbers.
2. Two numbers differ by 7. Three times the larger is five times the smaller olus 3. Find the length.
3. The length of a rectangle is 7cm more than twice the width. Find the dimensions if the perimeter is 62cm.

Answer by mananth(16946)

1. Two numebers have a sum of 26. Three times the first less twice the second is 3. Find the numbers.
1.x+y=26
2.3x-2y=3
multiply eq1 by 2 and add the equationa
2x+2y+3x-2y=52+3
5x=55
x=11
y=26-11=15
..
2. Two numbers differ by 7. Three times the larger is five times the smaller olus 3. Find the length.
1.x-y=7
3x=5y+3
2.3x-5y=3
multiply eq1 by 5 and add the equations
5x-5y+3x-5y=35+3
8x=38
x=4.75
y=-2.25
3. The length of a rectangle is 7cm more than twice the width. Find the dimensions if the perimeter is 62cm.
let width =x
length = x+7
Perimeter = 2L+2W
2x+2(x+7)
2x+2x+14=62
4x=62-14
4x=48
x=12cm the width
length = 12+7 =19cm